OpenStudy (anonymous):

How does the (255K)^4 get to (4.228 * 10^9) in this equation... E = (5.67 * 10-8 W m^-2 K^-4) * (255 K)^4 E = (5.67 * 10-8 W m^-2) * (4.228 * 10^9) E = 240 W m^-2

OpenStudy (anonymous):

255^4 = 4.228 * 10^9

OpenStudy (anonymous):

255^4 = 255*255*255*255 = 4228250625 = 4.228250625 × 10^9

OpenStudy (anonymous):

Okay thanks!